OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? a.three,seven,eleven b.four,eight,tweleve c.three,six,nine d.four,seven,eleven. ( Help me please! )

hartnn (hartnn):

simply put n= 1,2,3 so what is 4(1)-1 = ? 4(2)-1 = ? 4(3)-1 = ?

OpenStudy (anonymous):

From where is the sequence starting? If its starting from one the first term is 4*1-1. Second term is 4*2-1...and so on.

hartnn (hartnn):

@alexandria_15 did u get that ?

OpenStudy (anonymous):

Yes, Im doing all i can to understand and im kinda catching on. thank you,

hartnn (hartnn):

so did u get your answer ?